Thornton Corporation expects to incur indirect overhead costs of $171,825 per month and direct manufacturing costs of $12 per unit. The expected production activity for the first four months of the year are as follows.

January February March April
Estimated production in units 4,500 7,200 4,500 7,500

Required

  1. Calculate a predetermined overhead rate based on the number of units of product expected to be made during the first four months of the year.

  2. Allocate overhead costs to each month using the overhead rate computed in Requirement a.

  3. Calculate the total cost per unit for each month using the overhead allocated in Requirement b.
Answer #1

Answers

  • [a]

Jan

Feb

Mar

April

A

Monthly overhead

$171,825

$171,825

$171,825

$171,825

B = 4500 + 7200 + 4500 + 7500

Total units in 4 months

                     23,700

                     23,700

               23,700

               23,700

C = A/B

Predetremined Overhead rate per unit

$7.25

$7.25

$7.25

$7.25

  • [b]

Jan

Feb

Mar

April

A

Units

4500

7200

4500

7500

B

Predetremined Overhead rate per unit

$7.25

$7.25

$7.25

$7.25

C = A x B

Overhead cost allocated

$32,625

$52,200

$32,625

$54,375

  • [c]

Jan

Feb

Mar

April

A

Total overhead cost per unit

$7.25

$7.25

$7.25

$7.25

B

Manufacturing cost per unit

$12.00

$12.00

$12.00

$12.00

C = A+B

Total cost per unit

$19.25

$19.25

$19.25

$19.25
Answer #2

A. Allocation rate for overhead: 171,825 x 4 = $687,300

    Per unit: (171,825 x 4) / (4500 + 7200 + 4500 + 7500)

    = $29 per unit


B. Allocated Overhead Costs

    January (4500x29) = $130,500

    February = $208,800

    March =$ 130,500

    April = $217,500

    Total = $687,300


C. Total Cost Per Unit

January (184,500/4500)=$41.00


MonthJanFebMarApr
Num of Units           4,500          7,200              4,500              7,500
Exp Cost



     Overhead $   130,500 $  208,800 $      130,500 $      217,500
     Dir Cost $     54,000 $    86,400 $        54,000 $        90,000
Total Cost $   184,500 $  295,200 $      184,500 $      307,500
Cost per unit $       41.00 $      41.00 $           41.00 $           41.00

    

source: Completed the work myself
answered by: Nitch